OpenStudy (anonymous):

Divide 64 between 'A' and 'B' such that 3 times A's share is greater than 4 times B's share by 10.

OpenStudy (anonymous):

Solve the following for A and B\[\{A+B=64,3A=4B+10\}\]\[\{A=38,B=26\} \]\[3*38=4*26+10\text{ ?}\]\[114=114\]